# Looking through the Kerr disk

### v1

@@ I Introduction
The Kerr spacetime remains one of the most significant solutions of general relativity, both from physical [kw5g-d732] and mathematical [Andersson.2022, Klainerman.2022, Dafermos.2024] perspectives. Beyond its well-established role as the exterior geometry of rotating black holes [Kerr.1963], its maximal analytic extension reveals a remarkably rich global structure comprising two asymptotically flat regions connected through the ring singularity [carter_global_1968, Hawking.1973, chrusciel_geometry_2020]. Null geodesics that traverse this singular ring offer a rare glimpse into the less-explored interior of the Kerr geometry, providing valuable insight into the causal structure, optical properties, and potential physical interpretation of the extended manifold.
@@
By recasting the constants of motion in terms of impact parameters, we parametrize the observer’s celestial sphere and identify the subset corresponding to the inner throat [Helliwell.1975], a domain in which the radial potential remains positive and geodesics traverse the extended spacetime without reflection. We show that only photons with impact parameters inside this inner throat reach an observer at $r_{o}=$-$\infty$ from sources at $r_{s}=+\infty$. Analysing the angular potential, we find at most two distinct constant-$\theta$ geodesics within this region, one corresponding to the principal null direction, and the other existing only for specific combinations of the rotation parameter and observer inclination. Furthermore, our analysis reveals the presence of a forbidden range of polar angles, a region of the source’s sky that cannot be observed from the negative-$r$ domain.
@@ II The Kerr Metric
This section introduces the two coordinate systems for studying null geodesics in the Kerr spacetime [Kerr.1963]. Additionally, key features of the Kerr solution that are essential for the subsequent analysis are discussed.
We use Eddington-Finkelstein-like coordinates $(u,r,\theta,\psi)$, based on the Boyer-Lindquist coordinate system, tailored explicitly to the principal null congruences of the Kerr solution, i.e., adapted to in- and outgoing null geodesics [poisson_relativists_2004]. They are particularly useful for solving the geodesic equations (see Sec. III). Using these coordinates, the Kerr metric tensor is given by
@@
Three particular features of the Kerr spacetime are relevant to this work. The first is the event horizons, given by $\{\Delta=0\Leftrightarrow r=m\pm\sqrt{m^{2}-a^{2}}\}$. Horizons are present only for subcritical rotation ($0\leq a\leq m$); otherwise, the solution corresponds to a naked singularity. Note that both the Eddington-Finkelstein and the Kerr-Schild coordinates are regular across the horizons. A second feature is the ring singularity, defined by $\{\Sigma=0\Leftrightarrow r=0,\theta=\pi/2\}$, which is a curvature singularity [chrusciel_structure_2020]. The third is a causality-violating region [chrusciel_geometry_2020], also known as Carter’s time machine, characterized by regions where the Killing vector $\partial_{\psi}$ is timelike, i.e., $g_{\psi\psi}<0$. This region is generally not empty, and its name stems from the fact that a future-directed causal curve can connect any two points in this region.
@@ III Geodesic Equations of Motion in the Kerr Spacetime
The geodesic equations in Eddington-Finkelstein coordinates take the form
@@
where the dot represents differentiation with respect to some affine parameter $s$ along the geodesic and with the radial and angular potentials
@@
$\hskip 11.38109pt-\Delta\left(\mu^{2}r^{2}+Q+(L-aE)^{2}\right)\,$, (10)
$\Theta(\theta)$ $=Q-\cos^{2}\theta\left(a^{2}(\mu^{2}-E^{2})+\frac{L^{2}}{\sin^{2}\theta}\right)\,.$ (11)
Denoting the four-momentum by $p_{\alpha}=g_{\alpha\beta}\dot{x}^{\beta}$, the above equations contain the constants of motion corresponding to the energy $E=-p_{u}$, the angular momentum around the axis of symmetry $L=p_{\psi}$, the mass $\mu=\sqrt{-g_{\alpha\beta}\dot{x}^{\alpha}\dot{x}^{\beta}}$ of the test particle, and the redefined Carter’s constant [carter_global_1968]
$Q=p_{\theta}^{2}+\cos\theta\left(a^{2}\left(\mu^{2}-E^{2}\right)+\frac{L^{2}}{\sin^{2}\theta}\right)\,.$ (12)
For null geodesics ($\mu=0$), we can work with
@@
instead, meaning that only the sign of $E$ is important. We restrict the value of $E$ as follows: First, since we want to investigate geodesics that reach the region where $\partial_{u}$ is timelike, we can assume $E>0$ [oneill_geometry_2014]. Second, we can parameterize the geodesic so that the energy is normalized $E=1$, and thus $\lambda$ is the azimuthal angular momentum of the geodesic. Consequently, we will set $E=1$ in any numerical calculations.
In addition to the new constants of motion in (13), we take advantage of a cleverly chosen parameter $\tau$ along the geodesic, which makes it easier to solve the equations of motion for null geodesics. The so-called Mino time [mino_perturbative_2003] is defined by
@@ IV Visualization preliminaries
We want to study geodesics that start at some source at a positive radius $r_{s}>0$ and reach the observer located at $r_{o}=$-$\infty$. To this end, we integrate (15)-(18) back in time, starting from the observer at $r_{o}\rightarrow$-$\infty$ until we reach a specified source point at $r_{s}>0$.
@@ IV.1 Radial Motion
It should be clear that the radial motion for physical geodesics is restricted to the case where the radicand in (21) is nonnegative. Therefore, the real-valued roots of $R(r)$ determine the number of radial turning points of a null geodesic. For our study, we are interested in geodesics connecting a spacetime point with $r_{s}\gg 0$ to another spacetime point with $r_{o}\ll 0$. Such geodesics must consequently have zero radial turning points (cf. [oneill_geometry_2014]), i.e., $R(r)\neq 0$ for all $r$. Only if $\eta<0$, (21) can exhibit four complex roots and, therefore, no radial turning point [gralla_null_2020]. Geodesics with negative $\eta$ are also known as vortical null geodesics, and they are the main focus of our study.
@@ IV.1.1 Impact parameters
Consider an observer located at $(u_{o},r_{o},\theta_{o},\psi_{o})$ and a light source at $(u_{s},r_{s},\theta_{s},\psi_{s})$. [^1] Throughout this paper, the subscripts $o$ and $s$ indicate the coordinates corresponding to the spacetime location of an observer or a source, respectively. For $|r_{o}|\rightarrow\infty$, it is helpful to replace $\lambda$ and $\eta$ in (13) with the so-called impact parameters $\alpha$ and $\beta$, see [bardeen_timelike_1972]. Based on the assumption that an observer far away from the black hole measures the direction of photons that reach them relative to the center of symmetry of the spacetime (i.e., the observer’s line of sight is towards the center of the black hole), we define [^2] A detailed derivation of these impact parameters can be found in [bardeen_timelike_1972]. The only difference to (22) is an overall sign in the equation for $\alpha$. This stems from the fact that we consider $r_{o}<0$.
@@
Note that $\alpha$ and $\beta$ depend on the observer’s location; hence, the analysis of the radial potential becomes observer-dependent. Eq. (22) causes problems only at $\sin\theta_{o}=0$, corresponding to an observer located on the axis of symmetry. However, any geodesic going through the axis of symmetry (at which $\sin\theta=0$), and therefore especially those reaching an observer located in the axis of symmetry (with $\theta_{o}=0$), must necessarily have $\lambda=0$ (see, e.g., [oneill_geometry_2014]). The statement $\sin\theta_{o}=0\Leftrightarrow\lambda=0$ also becomes apparent when considering the inverses of (22):
@@ IV.1.2 Inner throat
As mentioned above, there are no turning points in the radial coordinate for geodesics within the inner throat, as these coincide with the real roots of $R(r)$. Thus, only null geodesics within the inner throat connect a light source at $r_{s}\gg 0$ and an observer at $r_{o}\ll 0$. Therefore, the inner throat serves as the field of view for our observer at $r_{o}\ll 0$ with a light source at $r_{s}\gg 0$. How the inner throat changes depending on the observer’s polar coordinate is depicted in Fig. 1. In particular, the left/right vertices, i.e., the inner throat boundary for $\beta=0$ and minimal/maximal $\alpha$, move to the right and become closer to each other when $\theta_{o}$ increases from 0 to $\pi/2$, cf. [Helliwell.1975].
@@ IV.2 Polar motion
A feature of vortical null geodesics ($\eta<0$) is that they are bound to one hemisphere, i.e., the geodesic is trapped either in the northern hemisphere $0\leq\theta<\pi/2$ or in the southern hemisphere $\pi/2<\theta\leq\pi$. The radicand in (25) is generally non-zero, so the geodesic exhibits oscillatory behavior in the polar coordinate. The polar turning points are given by [gralla_null_2020]
@@ IV.2.1 Geodesics with constant polar angle and principal null congruence
This does not have real-valued solutions except for $a_{crit,+}=0$ and $\theta_{o}=\pi/2$. Thus, $(\alpha_{+},\beta_{0})$ never crosses the inner throat boundary. Because for $a/m=0.99$ and $\theta_{o}=0$, the point $(\alpha_{+},\beta_{0})$ is inside the inner throat (see Fig. 1), we conclude that these geodesics are always inside the inner throat. Therefore, every observer at $r_{o}\ll 0$ with a polar angle $\theta_{o}\neq\pi/2$ receives photons with a constant polar angle at the impact parameters $(\alpha_{+},\,\beta_{0})$ originating at $r_{s}>0$. These geodesics correspond to the principal null congruence on which the Eddington-Finkelstein coordinates (1) are based. As such, they also have constant $\psi$ and $u$ (see [poisson_relativists_2004] or Appendix A).
@@ IV.2.2 Geodesic motion close to the axis of symmetry or the equatorial plane
On the other hand, from (26) and (27) together with (23), we find that inside the inner throat, the point closest to the equatorial plane (with $\theta_{+,\max}$) is reached at the right boundary of the throat, so for $(\alpha_{\max},\beta=0)$ [noauthor_httpswwwquantagonatmasters-thesis_nodate]. As $\theta_{+,\max}<\pi/2$ for $h=1$ and $\theta_{+,\max}>\pi/2$ for h=1h=-1, there is a “forbidden” region for the polar motion of vortical null geodesics inside the inner throat: they only reach values satisfying $0<\theta_{-}\leq\theta\leq\theta_{+}\leq\theta_{+,\max}<\pi/2$ for $h=1$ or $\pi/2<\theta_{+,\max}\leq\theta_{+}\leq\theta\leq\theta_{-}<\pi$ for h=1h=-1. Fig. 3 shows $\theta_{\pm}$ for the right vertex of the inner throat with $(\alpha_{\max},\beta=0)$ and $h=1$. For a given $\theta_{o}$, polar angles greater than $\theta_{+}$ constitute the forbidden polar angle region.
@@ V Solving the geodesic equations
Eq. (35) couples the $r$- and $\theta$-motion, and both integrals are equal to the elapsed Mino time $\tau$ along the geodesic. This approach is also taken in [gralla_null_2020], whose results we use whenever applicable. However, at some points, we use a different approach, which we indicate accordingly. The main difference in their calculations is that we work in Eddington-Finkelstein-like coordinates, which are globally regular.
@@
Lastly, a few words about the general approach of calculating the starting values of a null geodesic that reaches the observer with impact parameters $(\alpha,\beta)$. The fact that we fix the observer’s position and the impact parameters leads us to calculate the geodesic’s trajectory “back in time.” Thus, we calculate the trajectory of a vortical null geodesic originating from the observer at $r_{o}=$-$\infty$ and some $(\alpha,\beta)$ within the inner throat and terminating at the source point at $r_{s}>0$. This inversion of the geodesic direction is equivalent to exchanging the integration boundaries. Therefore, we can equivalently “integrate backward” by flipping the sign $\nu_{r}$. This is a standard approach in many calculations done for black hole visualizations (see, e.g., [james_gravitational_2015, gralla_lensing_2020]).
@@ V.1 Analytic solutions
To follow the general approach of [gralla_null_2020] for the calculation of the antiderivatives of (35)-(37), we need to define some additional quantities, which can be found in Appendix B. Included are formulas for the roots $r_{i}$ ($i=1,2,3,4$) of the radial potential $R(r)$, as well as for the quantities $a_{1}$, $a_{2}$, $b_{1}$, $b_{2}$, $C$, $D$, $g_{0}$, $k$, $k_{4}$, $x_{4}(r)$, $h$, and $X_{4}(\tau)$. We also define the integrals
@@ V.1.1 $\tau$-integral
The right-hand side equals $I_{0}$ in (39) and according to [gradshtein_table_2000] (§3.145), its antiderivative is
@@ V.1.2 $r$-integral
To obtain a formula for the radial component of a vortical null geodesic, the authors of [gralla_null_2020] inverted (43) by using basic properties of elliptic integrals and elliptic functions to derive a formula for $r_{o}(\tau)$ (see their Sec. 4 in Appendix B). The final formula is

### v2

@@ I Introduction
The Kerr spacetime remains one of the most significant solutions of general relativity, both from physical [19] and mathematical [2, 18, 9] perspectives. Beyond its well-established role as the exterior geometry of rotating black holes [17], its maximal analytic extension reveals a remarkably rich global structure comprising two asymptotically flat regions connected through the ring singularity [4, 13, 7]. Null geodesics that traverse this singular ring offer a rare glimpse into the less-explored interior of the Kerr geometry, providing valuable insight into the causal structure, optical properties, and potential physical interpretation of the extended manifold.
@@
By recasting the constants of motion in terms of impact parameters, we parametrize the observer’s celestial sphere and identify the subset corresponding to the inner throat [14], a domain in which the radial potential remains positive and geodesics traverse the extended spacetime without reflection. We show that only photons with impact parameters inside this inner throat reach an observer at $r_{o}=$-$\infty$ from sources at $r_{s}=+\infty$. Analysing the angular potential, we find at most two distinct constant-$\theta$ geodesics within this region, one corresponding to the principal null direction, and the other existing only for specific combinations of the rotation parameter and observer inclination. Furthermore, our analysis reveals the presence of a forbidden range of polar angles, a region of the source’s sky that cannot be observed from the negative-$r$ domain.
@@ II The Kerr Metric
This section introduces the two coordinate systems for studying null geodesics in the Kerr spacetime [17]. Additionally, key features of the Kerr solution that are essential for the subsequent analysis are discussed.
We use Eddington-Finkelstein-like coordinates $(u,r,\theta,\psi)$, based on the Boyer-Lindquist coordinate system, tailored explicitly to the principal null congruences of the Kerr solution, i.e., adapted to in- and outgoing null geodesics [22]. They are particularly useful for solving the geodesic equations (see Sec. III). Using these coordinates, the Kerr metric tensor is given by
@@
Three particular features of the Kerr spacetime are relevant to this work. The first is the event horizons, given by $\{\Delta=0\Leftrightarrow r=m\pm\sqrt{m^{2}-a^{2}}\}$. Horizons are present only for subcritical rotation ($0\leq a\leq m$); otherwise, the solution corresponds to a naked singularity. Note that both the Eddington-Finkelstein and the Kerr-Schild coordinates are regular across the horizons. A second feature is the ring singularity, defined by $\{\Sigma=0\Leftrightarrow r=0,\theta=\pi/2\}$, which is a curvature singularity [5]. The third is a causality-violating region [7], also known as Carter’s time machine, characterized by regions where the Killing vector $\partial_{\psi}$ is timelike, i.e., $g_{\psi\psi}<0$. This region is generally not empty, and its name stems from the fact that a future-directed causal curve can connect any two points in this region.
@@ III Geodesic Equations of Motion in the Kerr Spacetime
The geodesic equations in Eddington-Finkelstein coordinates for the wave vector $k^{\alpha}=(\dot{r},\dot{\theta},\dot{\phi},\dot{u})$, satisfying $k_{\alpha}k^{\alpha}=0$, take the form
@@
where $\nu_{r}=\pm 1$ and $\nu_{\theta}=\pm 1$ denote the signs of $\dot{r}$ and $\dot{\theta}$ respectively, and the dot denotes differentiation with respect to an affine parameter $s$ along the geodesic. The functions $R(r)$ and $\Theta(\theta)$ are the radial and angular potentials,
@@
$\hskip 11.38109pt-\Delta\left(Q+(L-aE)^{2}\right)\,$, (10)
$\Theta(\theta)$ $=Q-\cos^{2}\theta\left(-E^{2}a^{2}+\frac{L^{2}}{\sin^{2}\theta}\right)\,.$ (11)
The above equations contain the constants of motion corresponding to the energy $E=-k_{u}$, the angular momentum around the axis of symmetry $L=k_{\psi}$, and the redefined Carter’s constant [4]
$Q=k_{\theta}^{2}+\cos^{2}\theta\left(-E^{2}a^{2}+\frac{L^{2}}{\sin^{2}\theta}\right)\,.$ (12)
For null geodesics, we can work with
@@
instead, meaning that only the sign of $E$ is important. We restrict the value of $E$ as follows: First, since we want to investigate geodesics that reach the region where $\partial_{u}$ is timelike, we can assume $E>0$ [21]. Second, we can parameterize the geodesic so that the energy is normalized $E=1$, and thus $\lambda$ is the azimuthal angular momentum of the geodesic. Consequently, we will set $E=1$ in any numerical calculations.
In addition to the new constants of motion in (13), we take advantage of a cleverly chosen parameter $\tau$ along the geodesic, which makes it easier to solve the equations of motion for null geodesics. The so-called Mino time [20] is defined by
@@ IV Visualization preliminaries
We want to study geodesics that start at some source at a positive radius $r_{s}>0$ and reach the observer located at $r_{o}=$-$\infty$, hence we consider ingoing null geodesics. We want to emphasize that the light sources we consider throughout this paper are test sources that do not alter the spacetime or its symmetries. As such, they can be mathematically identified with points within the spacetime.
@@ IV.1 Radial Motion
It should be clear that the radial motion for physical geodesics is restricted to the case where the radicand in (21) is nonnegative. Therefore, the real-valued roots of $R(r)$ determine the number of radial turning points of a null geodesic. For our study, we are interested in geodesics connecting a spacetime point with $r_{s}\gg 0$ to another spacetime point with $r_{o}\ll 0$. Such geodesics must consequently have zero radial turning points (cf. [21]), i.e., $R(r)\neq 0$ for all $r$. Only if $\eta<0$, (21) can exhibit four complex roots and, therefore, no radial turning point [12]. Geodesics with negative $\eta$ are also known as vortical null geodesics, and they are the main focus of our study.
@@ IV.1.1 Impact parameters
Consider an observer located at $(u_{o},r_{o},\theta_{o},\psi_{o})$ and a light source at $(u_{s},r_{s},\theta_{s},\psi_{s})$. [^1] Throughout this paper, the subscripts $o$ and $s$ indicate the coordinates corresponding to the spacetime location of an observer or a source, respectively. For $|r_{o}|\rightarrow\infty$, it is helpful to replace $\lambda$ and $\eta$ in (13) with the so-called impact parameters $\alpha$ and $\beta$, see [3]. Based on the assumption that an observer far away from the black hole measures the direction of photons that reach them relative to the center of symmetry of the spacetime (i.e., the observer’s line of sight is towards the center of the black hole), we define [^2] A detailed derivation of these impact parameters can be found in [3]. The only difference to (22) is an overall sign in the equation for $\alpha$. This stems from the fact that we consider $r_{o}<0$.
@@
Note that $\alpha$ and $\beta$ depend on the observer’s location; hence, the analysis of the radial potential becomes observer-dependent. Eq. (22) causes problems only at $\sin\theta_{o}=0$, corresponding to an observer located on the axis of symmetry. However, any geodesic going through the axis of symmetry (at which $\sin\theta=0$), and therefore especially those reaching an observer located in the axis of symmetry (with $\theta_{o}=0$), must necessarily have $\lambda=0$ (see, e.g., [21]). The statement $\sin\theta_{o}=0\Leftrightarrow\lambda=0$ also becomes apparent when considering the inverses of (22):
@@ IV.1.2 Inner throat
As mentioned above, there are no turning points in the radial coordinate for geodesics within the inner throat, as these coincide with the real roots of $R(r)$. Thus, only null geodesics within the inner throat connect a light source at $r_{s}\gg 0$ and an observer at $r_{o}\ll 0$. Therefore, the inner throat serves as the field of view for our observer at $r_{o}\ll 0$ with a light source at $r_{s}\gg 0$. How the inner throat changes depending on the observer’s polar coordinate is depicted in Fig. 1. In particular, the left/right vertices, i.e., the inner throat boundary for $\beta=0$ and minimal/maximal $\alpha$, move to the right and become closer to each other when $\theta_{o}$ increases from 0 to $\pi/2$, cf. [14].
@@ IV.2 Polar motion
A feature of vortical null geodesics ($\eta<0$) is that they are bound to one hemisphere, i.e., the geodesic is trapped either in the northern hemisphere $0\leq\theta<\pi/2$ or in the southern hemisphere $\pi/2<\theta\leq\pi$. The radicand in (25) is generally non-zero, so the geodesic exhibits oscillatory behavior in the polar coordinate. The polar turning points are given by [12]
@@ IV.2.1 Geodesics with constant polar angle and principal null congruence
This does not have real-valued solutions except for $a_{crit,+}=0$ and $\theta_{o}=\pi/2$. Thus, $(\alpha_{+},\beta_{0})$ never crosses the inner throat boundary. Because for $a/m=0.99$ and $\theta_{o}=0$, the point $(\alpha_{+},\beta_{0})$ is inside the inner throat (see Fig. 1), we conclude that these geodesics are always inside the inner throat. Therefore, every observer at $r_{o}\ll 0$ with a polar angle $\theta_{o}\neq\pi/2$ receives photons with a constant polar angle at the impact parameters $(\alpha_{+},\,\beta_{0})$ originating at $r_{s}>0$. These geodesics correspond to the principal null congruence on which the Eddington-Finkelstein coordinates (1) are based. As such, they also have constant $\psi$ and $u$ (see [22] or Appendix A).
@@ IV.2.2 Geodesic motion close to the axis of symmetry or the equatorial plane
On the other hand, from (26) and (27) together with (23), we find that inside the inner throat, the point closest to the equatorial plane (with $\theta_{+,\max}$) is reached at the right boundary of the throat, so for $(\alpha_{\max},\beta=0)$ [23]. As $\theta_{+,\max}<\pi/2$ for $h=1$ and $\theta_{+,\max}>\pi/2$ for h=1h=-1, there is a “forbidden” region for the polar motion of vortical null geodesics inside the inner throat: they only reach values satisfying $0<\theta_{-}\leq\theta\leq\theta_{+}\leq\theta_{+,\max}<\pi/2$ for $h=1$ or $\pi/2<\theta_{+,\max}\leq\theta_{+}\leq\theta\leq\theta_{-}<\pi$ for h=1h=-1. Fig. 3 shows $\theta_{\pm}$ for the right vertex of the inner throat with $(\alpha_{\max},\beta=0)$ and $h=1$. For a given $\theta_{o}$, polar angles greater than $\theta_{+}$ constitute the forbidden polar angle region.
@@ V Solving the geodesic equations
Eq. (35) couples the $r$- and $\theta$-motion, and both integrals are equal to the elapsed Mino time $\tau$ along the geodesic. This approach is also taken in [12], whose results we use whenever applicable. However, at some points, we use a different approach, which we indicate accordingly. The main difference in their calculations is that we work in Eddington-Finkelstein-like coordinates, which are globally regular.
@@
Lastly, a few words about the general approach of calculating the starting values of a null geodesic that reaches the observer with impact parameters $(\alpha,\beta)$. The fact that we fix the observer’s position and the impact parameters leads us to calculate the geodesic’s trajectory “back in time.” Thus, we calculate the trajectory of a vortical null geodesic originating from the observer at $r_{o}=$-$\infty$ and some $(\alpha,\beta)$ within the inner throat and terminating at the source point at $r_{s}>0$. This technical inversion of the geodesic direction is equivalent to exchanging the integration boundaries. However, we emphasize that these geodesics still correspond to ingoing rays, crossing from the exterior region of the Kerr black hole to the interior. This is a standard approach in many calculations done for black hole visualizations (see, e.g., [16, 11]).
@@ V.1 Analytic solutions
To follow the general approach of [12] for the calculation of the antiderivatives of (35)-(37), we need to define some additional quantities, which can be found in Appendix B. Included are formulas for the roots $r_{i}$ ($i=1,2,3,4$) of the radial potential $R(r)$, as well as for the quantities $a_{1}$, $a_{2}$, $b_{1}$, $b_{2}$, $C$, $D$, $g_{0}$, $k$, $k_{4}$, $x_{4}(r)$, $h$, and $X_{4}(\tau)$. We also define the integrals
@@ V.1.1 $\tau$-integral
The right-hand side equals $I_{0}$ in (39) and according to [10] (§3.145), its antiderivative is
@@ V.1.2 $r$-integral
To obtain a formula for the radial component of a vortical null geodesic, the authors of [12] inverted (43) by using basic properties of elliptic integrals and elliptic functions to derive a formula for $r_{o}(\tau)$ (see their Sec. 4 in Appendix B). The final formula is